# HLSTester: Efficient Testing of Behavioral Discrepancies with LLMs for High-Level Synthesis

### v1

@@ 4. Experimental Results
Figure 13. Comparison of the average total time required to test all behavioral discrepancies.
Fig. 9 compares the proposed HLSTester framework with four baselines in terms of behavioral discrepancy detection speed. The x-axis denotes the elapsed time (i.e., the testing time budget), while the y-axis shows the average cumulative number of detected discrepancies. Each data point represents the average of 15 repeated simulations per task. By incorporating an LLM-driven reasoning chain, dynamic mutation, and redundancy filtering, the proposed HLSTester generates more effective test inputs to detect more behavioral discrepancies in the same time budget. In particular, the LLM-driven reasoning chain achieves an average of 1.81× speedup in detecting all behavioral discrepancies, while the dynamic mutation increases the speedup to an average of 2.28x, and the redundancy‑aware technique further accelerates the testing workflow by 15.73%.
To evaluate the total testing time required to detect all behavioral discrepancies, Fig. 13 compares the proposed HLSTester with the direct use of GPT and traditional fuzz testing. Each bar represents the average total testing time per benchmark, which includes HLS tool compilation, synthesis and simulation time, GPT interaction time, script execution time, and so on. As illustrated, the proposed HLSTester achieves an average speedup of 2.71$\times$ over the baseline, demonstrating the high efficiency achieved by optimizing the testing workflow.
5. Conclusion
In this paper, we propose HLSTester, an LLM-aided testing framework designed to efficiently detect behavioral discrepancies in HLS. Existing C/C++ testbenches are leveraged to guide the LLM in generating HLS-compatible testbenches that eliminate traditional C/C++ constructs incompatible with the HLS tool. Then, a backward slicing technique is employed to pinpoint the key variables, and a monitoring script is developed to record their runtime spectra, thereby enabling a detailed analysis of discrepancy symptoms. Moreover, a test input generation mechanism is proposed that leverages spectra feedback to guide dynamic mutation, which is incorporated with insights from an LLM-based progressive reasoning chain to increase the likelihood of revealing behavioral discrepancies. In addition, a redundancy-aware filtering technique is employed to skip repetitive simulations. Experimental results show that the proposed HLSTester significantly accelerates the testing workflow and achieves higher simulation pass rates compared with both the traditional method and the direct use of the LLM.

### v3

@@ IV Experimental Results
Figure 13: Comparison of the average total time required to test all discrepancies.
Fig. 9 compares the proposed HLSTester framework with four baselines in terms of behavioral discrepancy detection speed. The x-axis denotes the elapsed time (i.e., the testing time budget), while the y-axis shows the average cumulative number of detected discrepancies. By incorporating an LLM-driven reasoning chain, dynamic mutation, and redundancy filtering, the proposed HLSTester generates more effective test inputs to detect more behavioral discrepancies in the same time budget. In particular, the LLM-driven reasoning chain achieves an average of 1.81× speedup in detecting all behavioral discrepancies, while the dynamic mutation increases the speedup to an average of 2.28x, and the redundancy‑aware technique further accelerates the testing workflow by 15.73%.
To evaluate the total testing time required to detect all behavioral discrepancies, Fig. 13 compares the proposed HLSTester with the direct use of GPT and traditional fuzz testing. Each bar represents the average total testing time per benchmark, which includes HLS tool synthesis and simulation time, GPT interaction time, script execution time, and so on. As illustrated, the proposed HLSTester achieves an average speedup of 2.71$\times$ over the baseline, demonstrating the high efficiency achieved by optimizing the testing workflow.
V Conclusion
In this paper, we propose HLSTester, an LLM-aided testing framework designed to efficiently detect behavioral discrepancies in HLS. Existing C/C++ testbenches are used to guide the LLM in modifying them into HLS-compatible versions, with efficient test inputs generated by dynamic mutation with an LLM-based progressive reasoning chain to increase the likelihood of revealing behavioral discrepancies. A backward slicing technique is employed to pinpoint the key variables, and a monitoring script is developed to record their runtime spectra for detailed discrepancy analysis.Experimental results show that the proposed HLSTester significantly accelerates the testing workflow and achieves high simulation pass rates.
Acknowledgement
This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 504518248 and supported by TUM International Graduate School of Science and Engineering (IGSSE).